# NeuralRemaster: Phase-Preserving Diffusion for Structure-Aligned Generation

### v2

@@ Abstract
Standard diffusion corrupts data using Gaussian noise whose Fourier coefficients have random magnitudes and random phases. While effective for unconditional or text-to-image generation, corrupting phase components destroys spatial structure, making it ill-suited for tasks requiring geometric consistency, such as re-rendering, simulation enhancement, and image-to-image translation. We introduce Phase-Preserving Diffusion ($\phi$-PD), a model-agnostic reformulation of the diffusion process that preserves input phase while randomizing magnitude, enabling structure-aligned generation without architectural changes or additional parameters. We further propose Frequency-Selective Structured (FSS) noise, which provides continuous control over structural rigidity via a single frequency-cutoff parameter. $\phi$-PD adds no inference-time cost and is compatible with any diffusion model for images or videos. Across photorealistic and stylized re-rendering, as well as sim-to-real enhancement for driving planners, $\phi$-PD produces controllable, spatially aligned results. When applied to the CARLA simulator, $\phi$-PD improves CARLA-to-Waymo planner performance by 50%. The method is complementary to existing conditioning approaches and broadly applicable to image-to-image and video-to-video generation. Videos, additional examples, and code are available on our project page.
@@ 1 Introduction
Although these tasks are conceptually easier than generating an image from scratch, existing solutions are unnecessarily complex. Methods such as ControlNet [42], T2I-Adapter [21], and related variants attach auxiliary encoders, or adapter branches to inject structural input into the model. While effective, this introduces additional parameters and computational cost, paradoxically making structure-aligned generation harder than it should be.
We argue that this inefficiency stems not from the network architecture, but from the diffusion process itself. The forward diffusion process injects Gaussian noise, which destroys both the magnitude and phase components in the frequency domain. Classical signal processing [23, 30, 37], however, tells us that phase encodes structure while magnitude encodes texture. Destroying the phase means destroying the very spatial coherence that structure-aligned generation depends on, forcing the model to reconstruct structure from scratch.
@@
We evaluate $\phi$-PD across photorealistic re-rendering, stylized re-rendering and simulation enhancement for embodied-AI agents. $\phi$-PD consistently maintains geometry alignment while producing high-quality visual outputs, outperforming prior methods across both quantitative and qualitative metrics. When used to enhance CARLA simulations, $\phi$-PD improves planner transfer to the Waymo Open Dataset by 49%, substantially narrowing the sim-to-real gap. In summary, our contributions include:
@@ 2 Related Work
Diffusion Models. Diffusion models have become a dominant paradigm for generative modeling, capable of representing complex data distributions with remarkable fidelity [11, 16]. They progressively corrupt data into Gaussian noise through a forward diffusion process, then learn to invert this process via iterative denoising. This framework has demonstrated state-of-the-art performance across diverse domains, including image, video, and audio generation [31, 28, 3, 4, 12, 25, 17], as well as reinforcement learning [14, 36, 26] and robotics [1, 35, 5].
@@
Frequency-Domain Manipulation for Diffusion. Recent work has explored frequency domain operations for diffusion models. [7] argues that diffusion models of images perform approximate autoregression in the frequency domain. [40] shows that modifying the UNet frequency domain features significantly improves the generating quality for image or video generation. FreeDiff [38] introduces a fine-tuning free approach for image editing that employs progressive frequency truncation to refine the guidance of diffusion models. [27] proposed to use a frequency-dependent moving average during sampling. [2] proposes a training-free approach for image inpainting that optimizes the initial seed noise in the spectral domain. [9] proposes to generate optical illusion images using phase interpolation based on DDIM inversion. While these approaches modify the diffusion sampling process to achieve the desired behavior, $\phi$-PD introduces minimal changes to existing diffusion frameworks and preserves the original sampling dynamics.
@@
Structure-Aligned Generation with Diffusion. Most existing methods achieve structure-aligned generation by modifying the network architecture and introducing additional adaptation components. ControlNet [42] copies the entire U-Net encoder into a trainable encoder branch, which adds significant computation overhead. T2I-Adapter [21] reduces computation overhead using a lightweight adapter module but sacrifices control precision. Uni-ControlNet [43] enables simultaneous utilization of multiple local controls by training two adapters. OmniControl [32] integrates image conditions into Diffusion Transformer (DiT) architectures with only 0.1% additional parameters by re-using the VAE and transformer blocks of the base model. ControlNeXt [24] uses a lightweight convolutional module to inject control signals, and directly finetune selective parameters of the base model to reduce training costs and latency increase. SCEdit [15] proposes an efficient finetuning framework that edits skip connections using a lightweight module. NanoControl [13] aims to achieve efficient control with a LoRA-style control module. The above methods all rely on an additional module to incorporate the control signal, though some are more lightweight than others. CosmosTransfer [22] achieves multi-modal control by combining multiple ControlNet branches, demonstrating promising applications on physical tasks; however, multiple branches introduce significant computation overhead. In contrast, $\phi$-PD does not introduce any computation overhead or additional parameters while enabling universal spatial control.
@@
Training-Free Guidance Methods. Recently, several training-free methods have been developed. [41] introduced FreeDoM, which leverages off-the-shelf pre-trained networks to construct time-independent energy functions that guide generation. [6] proposed ZestGuide for zero-shot spatial layout conditioning, utilizing implicit segmentation maps extracted from cross-attention layers to align generation with input masks. [20] presented FreeControl, a training-free approach that enforces structure guidance with the base model feature extracted from the control signal. Although these methods avoid training cost, they introduce additional overhead at test time, either an external model, DDIM inversion, or multiple inferences of the base model. In contrast, $\phi$-PD can achieve training-free spatial control without any additional inference time overhead.
@@ 3.1 Frequency Domain Fundamentals
$I(x,y)$ $=\mathcal{F}^{-1}\{F(u,v)\}$
$=\sum_{u=0}^{W-1}\sum_{v=0}^{H-1}F(u,v)e^{2\pi j(ux/W+vy/H)}.$ (3)
Phase-Magnitude Separation in Signal Processing. Foundational work by Oppenheim et al. [23] shows that phase primarily determines spatial structure, while magnitude largely controls texture statistics. Mixing phase and magnitude from different images produces reconstructions whose spatial layout follows the source of the phase, not magnitude (see Figure 3). This observation motivates our approach: if diffusion destroys phase, it destroys spatial geometry; if we preserve phase, we preserve structure.
@@ 3.2 Phase-Preserving Diffusion
where the random magnitude $A_{\epsilon}$ can be either from the Fourier transforms of Gaussian noise:
$A_{\epsilon}=|\mathcal{F}\{\epsilon\}|,\quad\epsilon\sim\mathcal{N}(0,1),$ (7)
or sampled from Rayleigh distribution directly [10]:
$A_{\epsilon}=\sqrt{-2\ln U},\qquad U\sim\text{Uniform}(0,1).$ (8)
@@
Input $r=1$ $r=6$ $r=10$ $r=20$ $r=30$
Frequency Selective Structured (FSS) Noise.
In practice, we often want to control to what extent we keep the structure from the input image. Some tasks require strict structure preservation, while others benefit from partial freedom to reinterpret the scene. To provide this control, we introduce Frequency Selective Structured (FSS) noise, which only keep the image within a radius $r$ and use the phase from the noise for the remainder. We define a smooth frequency mask $M(u,v)$ based on the cutoff radius $r$:
@@
where $\sigma$ controls the smoothness of the transition.
@@
When the mask is all zero, we take the phase from Gaussian noise for all frequencies, then this FFS noise becomes Gaussian noise and $\phi$-PD becomes standard diffusion. Figure 4 visualizes FSS noise with different cutoff radius $r$. We can see that the noise becomes increasingly more structured with increasing $r$. Figure 5 shows images generated from the same input with different cutoff radius where the generated image aligns more tightly to the input with larger $r$.
@@ 3.3 Training Objective
The flow matching objective learns a vector field that transports the structured noise distribution to the target image distribution. During training, given a target image $I$ and a structured noise $\hat{\epsilon}$, and a timestep $t\in[0,1]$, an intermediate image $x_{t}$ is obtained using a linear combination between $I$ and $\hat{\epsilon}$ following Rectified Flows [19]:
@@
In the Fourier domain, the velocity becomes
@@
which has the same phase $\phi_{I}$ as the image; therefore, the trained model always generates images with the same phase as the input image, remaining structurally aligned by construction.
In DDPMs [11], data $x_{0}$ is gradually corrupted by Gaussian noise:
@@
The training objective remains identical,
$\mathcal{L}_{\text{$\phi$-PD{}}}=\mathbb{E}_{x_{0},\hat{\epsilon},t}\!\left[\|\hat{\epsilon}-\epsilon_{\theta}(x_{t},t)\|_{2}^{2}\right].$ (19)
The phase-preserving noise biases denoising toward structure-consistent trajectories, maintaining spatial alignment without altering the DDPM training objective.
@@ 3.4 Extension to Videos
$\phi$-PD can be directly repurposed for video generation by constructing phase-preserving noise frame-by-frame. For a video $\{I_{1},I_{2},...,I_{T}\}$, we construct structured noise for each frame, and concatenate them along the time dimension. We found that the best video generation results are when we first apply an image-based $\phi$-PD on the first frame, and then use the first-frame-conditioned video $\phi$-PD to generate the rest of the video. Similar to image generation, for video generation, our method requires no architectural changes either and simply supplies structured noise for each frame.
@@ 4 Experiments
We evaluate $\phi$-PD across three settings: photorealistic re-rendering, stylized re-rendering, and simulation enhancement for autonomous driving, comparing against state-of-the-art methods. Our goals are to (1) assess spatial alignment under appearance changes, (2) measure visual realism, and (3) quantify the impact on downstream embodied-AI tasks.
To demonstrate its broad applicability, we implement $\phi$-PD on three representative diffusion models: SD 1.5, FLUX-dev, and WAN 2.2 14B, which vary in size, formulation, and modality, covering both image and video generation.
@@ 4.1.1 Datasets
UnrealCV [^1] https://github.com/unrealcv/unrealcv is a open-source tool that includes multiple assets. We created a diverse test set consisting of 5,000 images across all available assets, for a total of around 200 scenes. Figure 6 shows examples from this test set. This dataset covers a diverse range of scenes, including outdoor and indoor, city and natural etc, with geometry diversity while lacking photorealism. This dataset evaluates photorealistic enhancement and structure preservation.
ImageNetR is a test set proposed by [34], including 29 images of various objects and styles. While the original dataset provides prompts, these are generic image editing prompts. Since our work primarily focuses on re-rendering, we keep the editing prompts with style hints in the original dataset and added additional style prompts, resulting in a total of 8 prompts for each image. This benchmark assesses stylized re-rendering and structure preservation.
CARLA is an open-source driving simulator [8]. We collect 5.5 hours of driving videos from CARLA Town 4 using the simulator’s default autopilot. We then split the videos into 25 second clips and annotate a caption for each clip. For simulation enhancement, we use these captions combined with the style hint “A photorealistic video of driving”. We evaluate the effectiveness of sim-to-real transfer by testing the CARLA-trained planner on Waymo’s WOD-E2E [39] validation set.
@@ 4.1.2 Model architecture
We integrate $\phi$-PD into:
  • SD 1.5 (image DDPM)
  • FLUX-dev (image flow matching)
  • Wan2.2-14B (video flow matching)
@@
We either fully finetune or LoRA-finetune each model using phase-preserving noise; no architectural changes are introduced. Notably, this finetuning is highly efficient: adapting the Wan2.2-14B video model with LoRA required only a single GPU while still yielding high-quality results, further demonstrating the lightweight nature of $\phi$-PD. Please refer to the Appendix for additional implementation details and ablation studies.
4.1.3 Evaluation Metrics
For photorealistic re-rendering, we use our UnrealCV dataset and apply different methods to make images photorealistic. We use the style prompt ”A high-quality picture” combined with the caption extracted from the images using ChatGPT.
Visual quality. We define an appearance score (AS) that measures how successful re-rendering is by looking at the ratio of CLIP similarity between positive and negative prompts:
$AS=\frac{x^{\top}t_{p}}{x^{\top}t_{n}}$ (20)
where $x$ indicates the CLIP embedding of the re-rendered image; $t_{n}$ the CLIP embedding of the negative prompt; and $t_{p}$ the CLIP embedding of the positive prompt. For the positive prompt we use “Photo, camera captured, picture, photorealistic”, and for the negative prompt we use “Game, render, cartoon, unreal”.
Structural alignment. Besides visual quality, successful re-rendering also requires preserving the structure from the original image. To evaluate the structural alignment, we compute the error between the depth map of the original image and the generated image, using metrics SSIM (Structural Similarity Index Measure) and ABSREL (Absolute Relative error).
For stylized re-rendering, we evaluate in a similar way as for photorealistic re-rendering. We compute the appearance score with the target style prompt as positive and the original style hint of the original image as negative. We also use the error between depth maps to evaluate structural alignment.
For simulation enhancement, we train an end2end planner for each compared method on the corresponding re-rendered image. We then test the planner’s trajectories on Waymo’s WOD-E2E validation set using distance-based metrics from the demonstrated driving, including Average Displacement Error (ADE) and Final Displacement Error (FDE).
4.2 Results
Table 1: Quantitative evaluation results for photorealistic re-rendering on UnreadlCV.
Model Input images ControlNet-Tile SDEdit Ours
AS 0.9485 0.9733 0.9782 1.0008
SSIM - 0.8781 0.8883 0.8982
ABSREL - 0.5936 0.4938 0.4569
Table 2: Quantitative evaluation for stylized re-rendering.
Model ControlNet-Tile SDEdit PNP Ours
AS 1.3167 1.4243 1.4726 1.4709
SSIM 0.8831 0.7638 0.8498 0.8502
ABSREL 0.6684 1.0336 0.8194 0.7949
Photorealistic Re-Rendering. Quantitative results on UnrealCV re-rendering are summarized in Table 1, where all methods are implemented using the SD 1.5-based models for fair quantitative comparison. Qualitative examples are shown in Figure 6, which compares our Flux-based model against stronger recent models such as FLUX-Kontext, and Qwen-Edit. Across both settings, all methods improve photorealism-as reflected by higher AS scores than the input images while $\phi$-PD achieves the highest photorealism and superior structure alignment. We observe that QWen-Edit produces visually high-quality results but often fails to maintain structural alignment with the input image; for example, in the first four cases, it enlarges the main subjects significantly. FLUX-Kontext aligns better with the input structure but provides only limited improvement in visual quality. Our method achieves both high visual fidelity and consistent structural alignment across frames.
Stylized Re-rendering. Results of stylized re-rendering are presented in Table 2, with representative examples shown in Figure 8. All models are based on SD 1.5. This task evaluates the model’s ability to alter appearance while preserving scene structure. As shown, $\phi$-PD produces visually coherent stylizations that maintain object boundaries and spatial consistency, while prior methods often distort geometry or introduce texture misalignment. Quantitatively, $\phi$-PD achieves a better trade-off between style strength and structural fidelity.
Pencil Sketch of a Castle. ”
Picture of a Husky”
  Input PnP SDEdit ControlNet Ours
Figure 8: Results on stylized rerendering.
Figure 9: Video re-rendering results for simulation enhancement. From top to bottom are: input, Cosmos-Transfer2.5 [22], ours.
5 Conclusion
We introduced Phase-Preserving Diffusion ($\phi$-PD), a simple yet effective reformulation of the diffusion process that replaces Gaussian noise with structured noise that preserves image phase while randomizing the magnitude in the frequency domain. This simple change retains spatial alignment throughout sampling without modifying the architecture, altering training objectives, or introducing inference-time overhead. We also introduced Frequency-Selective Structured (FSS) noise, which provides continuous control over structural alignment rigidity through a single frequency cutoff parameter, making it broadly applicable to different applications. Across photorealistic re-rendering, stylized re-rendering, and simulation enhancement for driving, $\phi$-PD demonstrates strong spatial fidelity and visual realism. When applied to CARLA re-rendering, $\phi$-PD significantly improves zero-shot planner transfer to the Waymo dataset, narrowing the sim-to-real gap.
Appendix A Additional Implementation Details
We implemented $\phi$-PD based on three models: SD 1.5 [29], FLUX-dev [18] and Wan2.2-14b [33]. We use the implementation of these models from DiffSynthStudio [^2] https://github.com/modelscope/DiffSynth-Studio This section describes the additional implementation details for the experiment with each model.
A.1 Training and Inference Details

### v3

@@ Abstract
Standard diffusion corrupts data using Gaussian noise whose Fourier coefficients have random magnitudes and random phases. While effective for unconditional or text-to-image generation, corrupting phase components destroys spatial structure, making it ill-suited for tasks requiring geometric consistency, such as re-rendering, simulation enhancement, and image-to-image translation. We introduce Phase-Preserving Diffusion ($\phi$-PD), a model-agnostic reformulation of the diffusion process that preserves input phase while randomizing magnitude, enabling structure-aligned generation without architectural changes or additional parameters. We further propose Frequency-Selective Structured (FSS) noise, which provides continuous control over structural rigidity via a single frequency-cutoff parameter. $\phi$-PD adds no inference-time cost and is compatible with any diffusion model for images or videos. Across photorealistic and stylized re-rendering, as well as sim-to-real enhancement for driving planners, $\phi$-PD produces controllable, spatially aligned results. When applied to the CARLA simulator, $\phi$-PD significantly improves sim-to-real planner transfer performance. The method is complementary to existing conditioning approaches and broadly applicable to image-to-image and video-to-video generation. Videos, additional examples, and code are available on our project page.
@@ 1 Introduction
Although these tasks are conceptually easier than generating from scratch, existing solutions are unnecessarily complex. Methods such as ControlNet [zhang2023adding], T2I-Adapter [mou2024t2i], and related variants attach auxiliary branches to inject structural input into the model. While effective, this introduces additional parameters and computational cost, paradoxically making structure-aligned generation harder than it should be.
We argue that this inefficiency stems not from the network architecture, but from the diffusion process itself. The forward diffusion process injects Gaussian noise, which destroys both the magnitude and phase components in the frequency domain. Classical signal processing [oppenheim1981importance, ruderman1994statistics, wang2005translation], however, tells us that phase encodes structure while magnitude encodes texture. Destroying the phase means destroying the very spatial coherence that structure-aligned generation depends on, forcing the model to reconstruct structure from scratch.
@@
We evaluate $\phi$-PD across photorealistic re-rendering, stylized re-rendering and simulation enhancement for embodied-AI agents. $\phi$-PD consistently maintains geometry alignment while producing high-quality visual outputs, outperforming prior methods across both quantitative and qualitative metrics. When used to enhance CARLA simulations, $\phi$-PD improves planner transfer to the Waymo Open Dataset by 50%, substantially narrowing the sim-to-real gap. In summary, our contributions include:
@@ 2 Related Work
Diffusion Models. Diffusion models have become a dominant paradigm for generative modeling, capable of representing complex data distributions with remarkable fidelity [ho2020denoising, karras2022elucidating]. They progressively corrupt data into Gaussian noise through a forward diffusion process, then learn to invert this process via iterative denoising. This framework has demonstrated state-of-the-art performance across diverse domains, including image, video, and audio generation [saharia2022photorealistic, ramesh2022hierarchical, balaji2022ediff, chen2023videocrafter1, ho2022video, popov2021grad, kong2020diffwave], as well as reinforcement learning [janner2022planning, wang2022diffusion, psenka2023learning] and robotics [ajay2022conditional, urain2023se, chi2023diffusion].
@@
Frequency-Domain Manipulation for Diffusion. Recent work has explored frequency domain operations for diffusion models. [dielemanblog] argues that diffusion models of images perform approximate autoregression in the frequency domain. [falck2025fourier] shows that the standard diffusion forward process corrupts high-frequency components faster than low-frequency ones, and introduces an alternate process that corrupts all frequencies at the same rate. [yu2025dmfft] shows that modifying the UNet frequency domain features significantly improves the generating quality for image or video generation. FreeDiff [wu2024freediff] introduces a fine-tuning free approach for image editing that employs progressive frequency truncation to refine the guidance of diffusion models. [zhang2025training] proposes a training-free style transfer method that modulates the intermediate samples with the image phase. [qian2024boosting] proposed to use a frequency-dependent moving average during sampling. [baek2025sonic] proposes a training-free approach for image inpainting that optimizes the initial seed noise in the spectral domain.
@@
Structure-Aligned Generation with Diffusion. Most existing methods achieve structure-aligned generation by modifying the network architecture and introducing additional adaptation components. ControlNet [zhang2023adding] copies the entire U-Net encoder into a trainable encoder branch, which adds significant computation overhead. T2I-Adapter [mou2024t2i] reduces computation overhead using a lightweight adapter module but sacrifices control precision. Uni-ControlNet [zhao2023unicontrolnet] enables simultaneous utilization of multiple local controls by training two adapters. OmniControl [tan2024ominicontrol2] integrates image conditions into Diffusion Transformer (DiT) architectures with only 0.1% additional parameters by re-using the VAE and transformer blocks of the base model. ControlNeXt [peng2024controlnext] uses a lightweight convolutional module to inject control signals, and directly finetune selective parameters of the base model to reduce training costs and latency increase. SCEdit [jiang2024scedit] proposes an efficient finetuning framework that edits skip connections using a lightweight module. NanoControl [huang2024nanocontrol] aims to achieve efficient control with a LoRA-style control module. The above methods all rely on an additional module to incorporate the control signal, though some are more lightweight than others. CosmosTransfer [Alhaija2025CosmosTransfer1CW] achieves multi-modal control by combining multiple ControlNet branches, demonstrating promising applications on physical tasks; however, multiple branches introduce significant computation overhead. In contrast, $\phi$-PD does not introduce any computation overhead or additional parameters while enabling universal spatial control.
@@
Training-Free Guidance Methods. Recently, several training-free methods have been developed. [yu2023freedom] introduced FreeDoM, which leverages off-the-shelf pre-trained networks to construct time-independent energy functions that guide generation. [couairon2023zeroshot] proposed ZestGuide for zero-shot spatial layout conditioning, utilizing implicit segmentation maps extracted from cross-attention layers to align generation with input masks. [mo2024freecontrol] presented FreeControl, a training-free approach that enforces structure guidance with the base model feature extracted from the control signal. Although these methods avoid training cost, they introduce additional overhead at test time, either an external model, DDIM inversion, or multiple inferences of the base model. In contrast, $\phi$-PD achieves spatial control without any additional inference time overhead.
@@ 3.1 Frequency Domain Fundamentals
$I(x,y)$ $=\mathcal{F}^{-1}\{F(u,v)\}=\sum_{u=0}^{W-1}\sum_{v=0}^{H-1}F(u,v)e^{2\pi j(ux/W+vy/H)}.$ (3)
Phase-Magnitude Separation in Signal Processing. Foundational work by Oppenheim [oppenheim1981importance] shows that phase primarily determines spatial structure, while magnitude largely controls texture statistics. Mixing phase and magnitude from different images produces reconstructions whose spatial layout follows the source of the phase, not magnitude (see Figure 3). This observation motivates our approach: if diffusion destroys phase, it destroys spatial geometry; if we preserve phase, we preserve structure.
@@ 3.2 Phase-Preserving Diffusion
where the random magnitude $A_{\epsilon}$ can be obtained from the Fourier transform of Gaussian noise:
$A_{\epsilon}=|\mathcal{F}\{\epsilon\}|,\quad\epsilon\sim\mathcal{N}(0,I).$ (7)
Alternatively, $A_{\epsilon}$ can be sampled from a scaled Rayleigh distribution [GoodmanStatisticalOptics]:
$A_{\epsilon}=\sqrt{-N\ln U},\quad U\sim\text{Uniform}(0,1),$ (8)
where $N=H\times W$, matching the magnitude statistics of the DFT of unit-variance Gaussian noise.
@@
Frequency Selective Structured (FSS) Noise. In practice, we often want to control to what extent we keep the structure from the input image. Some tasks require strict structure preservation, while others benefit from partial freedom to reinterpret the scene. To provide this control, we introduce Frequency Selective Structured (FSS) noise, which only keep the image within a radius $r$ and use the phase from the noise for the remainder. We define a smooth frequency mask $M(u,v)$ based on the cutoff radius $r$:
@@
where $(u,v)$ are centered frequency coordinates with the DC component at the origin, and $r$ is specified in pixels relative to the frequency grid of size $W\times H$. In practice, we use fftshift to center the spectrum before applying the mask.
@@
When the mask is all zero, we take a random phase for all frequencies, then this FFS noise becomes Gaussian noise and $\phi$-PD becomes standard diffusion. Figure 5 visualizes FSS noise with different cutoff radius $r$. We can see that the noise becomes increasingly more structured with increasing $r$. Figure 5 shows images generated from the same input with different cutoff radius where the generated image aligns more tightly to the input with larger $r$.
Input $r=1$ $r=6$ $r=10$ $r=20$ $r=30$
@@ 3.3 Training Objective
The flow matching objective learns a vector field that transports the structured noise distribution to the target image distribution. During training, given a target image $I$ and a structured noise $\hat{\epsilon}$, and a timestep $t\in[0,1]$, an intermediate image $x_{t}$ is obtained using a linear combination between $I$ and $\hat{\epsilon}$ following Rectified Flows [liu2022flow]:
@@
In the Fourier domain, when using full phase-preservation, the velocity becomes
@@
which has phase $\phi_{I}$ or $\phi_{I}+\pi$, preserving structural alignment. With FSS noise, phase is preserved within the cutoff radius $r$, providing controllable alignment strength.
In DDPMs [ho2020denoising], data $x_{0}$ is gradually corrupted by Gaussian noise:
@@
The training objective follows the same form,
$\mathcal{L}_{\text{$\phi$-PD{}}}=\mathbb{E}_{x_{0},\hat{\epsilon},t}\!\left[\|\hat{\epsilon}-\epsilon_{\theta}(x_{t},t)\|_{2}^{2}\right],$ (19)
but the expectation is now over structured noise $\hat{\epsilon}$ rather than Gaussian noise $\epsilon$. Although $\hat{\epsilon}$ is non-Gaussian and formally violates standard DDPM assumptions, we observe high-quality, structure-aligned outputs using standard samplers after finetuning with this objective.
@@ 3.4 Extension to Videos
$\phi$-PD extends to video by constructing phase-preserving noise frame-by-frame. For a video $\{I_{1},I_{2},...,I_{T}\}$, we construct structured noise for each frame and concatenate along the time dimension. Since current video diffusion models produce lower-fidelity individual frames than image models, we adopt a two-stage pipeline: generating the initial frame with image-based $\phi$-PD, then extending temporally with first-frame-conditioned video $\phi$-PD. No architectural changes are required.
@@ 4 Experiments
Pencil Sketch of a Castle. "
Picture of a Husky"
  Input PNP SDEdit ControlNet Ours
Figure 7: Stylized re-rendering results.
We evaluate $\phi$-PD across three settings: photorealistic re-rendering, stylized re-rendering, and simulation enhancement for autonomous driving, comparing against state-of-the-art methods. To demonstrate its broad applicability, we implement $\phi$-PD on three representative diffusion models: SD 1.5, FLUX-dev, and WAN 2.2 14B, which vary in size, formulation, and modality, covering both image and video generation. Please refer to the supplementary materials for additional results and code.
During training, we use real images: construct structured noise from the image’s phase, corrupt the image, and train the model to denoise. At inference, $\phi$-PD supports two inference modes. (1) Structured SDEdit: add structured noise to the source image at some noise level $t$, then denoise. (2) From noise: denoise from pure structured noise without any pixel contribution from the source; structure is encoded entirely in the phase. Mode (1) generally yields stronger structural alignment; mode (2) offers more flexibility with larger appearance changes.
@@ 4.1.1 Datasets
UnrealCV [^1] https://github.com/unrealcv/unrealcv is a open-source tool that includes multiple assets. We created a diverse test set consisting of around 5,000 images across all available assets, for a total of around 200 scenes. Figure 6 shows examples from this test set. This dataset covers a diverse range of scenes, including outdoor and indoor, city and natural etc, with geometry diversity while lacking photorealism. This dataset evaluates photorealistic enhancement and structure preservation.
ImageNetR is a test set proposed by [pnp], including 29 images of various objects and styles. While the original dataset provides prompts, these are generic image editing prompts. Since our work primarily focuses on re-rendering, we keep the editing prompts with style hints in the original dataset and added additional style prompts, resulting in a total of 8 prompts for each image. This benchmark assesses stylized re-rendering and structure preservation.
CARLA is an open-source driving simulator [dosovitskiy2017carla]. We collect 5.5 hours of driving videos from CARLA Town 4 using the simulator’s default autopilot. We then split the videos into 25 second clips and annotate a caption for each clip. For simulation enhancement, we use these captions combined with the style hint “A photorealistic video of driving". We evaluate the effectiveness of sim-to-real transfer by testing the CARLA-trained planner on Waymo’s WOD-E2E [xu2025wode2e] validation set.
@@ 4.1.2 Model architecture
We integrate $\phi$-PD into: SD 1.5 (image DDPM), FLUX-dev (image flow matching), and Wan2.2-14B (video flow matching). We either fully finetune or LoRA-finetune each model using phase-preserving noise; no architectural changes are introduced. Notably, this finetuning is highly efficient: adapting the Wan2.2-14B video model with LoRA required only a single GPU while still yielding high-quality results, further demonstrating the lightweight nature of $\phi$-PD. To evaluate structure alignment, we compute LPIPS [zhang2018unreasonable] between the input and output pairs, as well as the error between their depth map using metrics SSIM (Structural Similarity Index Measure) and ABSREL (Absolute Relative error). To evaluate text prompt alignment, we compute CLIP [radford2021learning] similarity between the generated images and the input text prompt.
4.1.3 Training and Inference Details